# ProRefine: Inference-time Prompt Refinement with Textual Feedback

### v1

@@ 4.3 Results
ProRefine improves performance by $3-16$ percentage points over CoT, with significant gains observed for Llama3.1-8B-instruct. It outperforms TextGrad on 2 out of 3 models, yielding a $2-3$ percentage point advantage. However, a performance drop of $2$ points is observed for Llama3.2-1B-instruct.
@@
Performance gains over CoT range from $8-37$ percentage points, with significant improvements for Llama3.2-3B-instruct and Llama3.1-8B-instruct. ProRefine surpasses TextGrad on 1 of 3 models with a $2$-point gain, but performance drops of $11-14$ points are observed for Llama3.2-1B-instruct and Llama3.2-3B-instruct.
@@
ProRefine achieves $2.4-20.4$ percentage points improvement over CoT, with significant improvement observed for all the models; however, a slight performance drop ($1.2$) is observed for Llama3.2-3B-instruct. It outperforms TextGrad on all models, achieving a $2.1-19.1$ percentage point gain with significant results observed for Llama3.2-1B-instruct and Llama3.2-3B-instruct models. Minor performance drop of $0.4-2.1$ is observed for Llama3.2-3B-instruct and Llama3.1-8B-instruct.
@@
Performance improves by $0.9-11.9$ percentage points over CoT, with significant gains for Llama3.2-1B-instruct and Llama3.1-8B-instruct. ProRefine outperforms TextGrad across all models, with $1.7-12.4$ percentage point gains and significant results for Llama3.2-1B-instruct.
@@
Gains over CoT range from $0.9-2$ percentage points, but declines of $5.4-22$ points are also observed. ProRefine exceeds TextGrad on 2 of 3 models, with $2.7-4.7$ percentage point gains, though performance drops of $10.6-18.1$ points are also recorded.
Our results demonstrate that using ProRefine with an optimal verifier significantly improves performance across all tasks, achieving the best results in 13 out of 15 cases, highlighting the critical role of verifier quality. Notably, the number of significant improvements increases with larger model sizes.
We also observe that ProRefine can improve the performance of lesser capable models Llama3.2-3B-instruct and Llama3.1-8B-instruct to that of more capable ones Llama3.1-8B-instruct and Llama3.1-70B-instruct respectively. In our experiments using Llama3.1-70B-instruct for $LLM_{task}$, some experiments yielded an accuracy of 1, suggesting potential overfitting. Additionally, we experimented using ProRefine with a comparatively smaller model (Llama3.1-8B-instruct) for $LLM_{feedback}$ and $LLM_{optimizer}$. However, this configuration did not consistently yield substantial improvements and, in some cases, even degraded performance. This aligns with prior findings indicating that smaller models are less capable of providing high-quality feedback [31, 2].
Figure 4: Average number of prompt refinement iterations.
5 Discussion
Regarding RQ1, the results demonstrate that ProRefine is a broadly applicable method utilizing textual feedback for improving LLM performance at inference-time. The “performance gap bridging” effect is particularly noteworthy, suggesting that ProRefine can be an effective alternative to simply scaling up model size, potentially avoiding costly fine-tuning, which could prove critical in resource-constrained settings.
@@ 5 Discussion
Regarding RQ3, the results highlight that employing a high-quality verifier is crucial for significantly improving task performance at inference time. Beyond enhancing performance, the verifier also reduces computational cost during inference by guiding the refinement process. Moreover, it opens up promising avenues for future work, where an optimizer could be designed to maximize rewards guided by the verifier’s assessments.
ProRefine can offer a degree of interpretability by exposing the outputs from $LLM_{feedback}$, allowing insights into the model’s reasoning process. Figures 2 and 5 demonstrate cases where model feedback successfully improves the output and where it fails, respectively. Although evaluated on reasoning and math tasks, ProRefine is general and applicable to other tasks.
While ProRefine improves accuracy, it does increase the computational cost during inference. Each step requires calls to both the task LLM and the feedback/optimization LLMs. The trade-off between accuracy and computational cost should be considered when deploying ProRefine for test-time scaling. However, the cost is still considerably lower than full model retraining or fine-tuning. Figure 4 shows the average number of prompt refinement iterations using our standard approach, ProRefine (Verifier), for all datasets and models. Even though we use the same $n=25$ for all datasets and models, Figure 4 suggests that the granularity of the feedback ($k$) and the length of optimization ($n$) may be tuned for specific tasks to achieve an optimal balance between performance and computational cost.
Figure 5: ProRefine example. Given an input query and an initial prompt, the task model ($LLM_{task}$) gives an incorrect answer. Prorefine uses two additional models, $LLM_{feedback}$ and $LLM_{optimizer}$, to iteratively improve the prompt as the $LLM_{task}$ generates its response. Refining the prompt during generation allows the feedback model to target local regions of the response, providing finer-grained feedback than waiting for the response to complete. However, the optimizer is unable to utilize the feedback in this case.
6 Limitations and Future Work
@@
  • Computational Cost and Scalability: ProRefine’s iterative process increases computational cost compared to single-shot prompting. Scaling to substantially larger or more complex tasks presents both computational and methodological challenges.
  • Generalizability: Evaluation is currently limited to reasoning and mathematical tasks. Further research is needed to assess performance across a broader range of NLP tasks and domains.
  • Model Dependence and Feedback Quality: ProRefine relies on a capable LLM for feedback, which may limit accessibility. The quality and potential biases of this feedback are crucial factors impacting performance and require further investigation.
  • Evaluation and Bias: Using an LLM for evaluation introduces potential biases. More comprehensive human evaluations and robust methods for mitigating evaluator bias are needed.
  • Hyperparameter and Feedback Loop: Current approach requires manual hyperparameter tuning, and iterative prompting might cause issues, and automatic ways need further research.
7 Conclusion
We introduced ProRefine, a novel, practical, and inference-time prompt optimization method for agentic workflows. ProRefine leverages LLM-generated textual feedback to dynamically refine prompts, leading to significant performance improvements without requiring additional training or ground-truth labels. The results across diverse datasets demonstrate ProRefine’s effectiveness for multi-step reasoning tasks and its ability to bridge the performance gap between smaller and larger LLMs. The inference-time nature makes ProRefine readily deployable and adaptable, contributing to more efficient, transparent, and accessible AI systems.

### v2

@@ Results
ProRefine improves performance by 3163-16 percentage points over CoT, with significant gains observed for Llama3.1-8B-instruct. It outperforms TextGrad on 2 out of 3 models, yielding a 232-3 percentage point advantage. However, a performance drop of $2$ points is observed for Llama3.2-1B-instruct.
@@
Performance gains over CoT range from 8378-37 percentage points, with significant improvements for Llama3.2-3B-instruct and Llama3.1-8B-instruct. ProRefine surpasses TextGrad on 1 of 3 models with a $2$-point gain, but performance drops of 111411-14 points are observed for Llama3.2-1B-instruct and Llama3.2-3B-instruct.
@@
ProRefine achieves 2.420.42.4-20.4 percentage points improvement over CoT, with significant improvement observed for all the models; however, a slight performance drop ($1.2$) is observed for Llama3.2-3B-instruct. It outperforms TextGrad on all models, achieving a 2.119.12.1-19.1 percentage point gain with significant results observed for Llama3.2-1B-instruct and Llama3.2-3B-instruct models. Minor performance drop of 0.42.10.4-2.1 is observed for Llama3.2-3B-instruct and Llama3.1-8B-instruct.
@@
Performance improves by 0.911.90.9-11.9 percentage points over CoT, with significant gains for Llama3.2-1B-instruct and Llama3.1-8B-instruct. ProRefine outperforms TextGrad across all models, with 1.712.41.7-12.4 percentage point gains and significant results for Llama3.2-1B-instruct.
@@
Gains over CoT range from 0.920.9-2 percentage points, but declines of 5.4225.4-22 points are also observed. ProRefine exceeds TextGrad on 2 of 3 models, with 2.74.72.7-4.7 percentage point gains, though performance drops of 10.618.110.6-18.1 points are also recorded.
Our results demonstrate that using ProRefine with an optimal verifier significantly improves performance for all tasks, achieving the best results in 13 out of 15 cases, highlighting the critical role of verifier quality. Notably, number of significant improvements increases with larger model sizes.
We also observe that ProRefine enables smaller models, such as Llama3.2-3B-instruct and Llama3.1-8B-instruct, to approach the zero-shot performance of larger models like Llama3.1-8B-instruct and Llama3.1-70B-instruct, respectively. In our experiments using Llama3.1-70B-instruct for $LLM_{task}$, some experiments yielded an accuracy of 1, suggesting potential data contamination. We also experimented using the full chat history at each step, but it failed to produce consistent improvements over our stateless approach. Additionally, we experimented using ProRefine with a comparatively smaller model (Llama3.1-8B-instruct) for $LLM_{feedback}$ and $LLM_{optimizer}$. However, this configuration also failed to consistently yield substantial improvements and, in some cases, even degraded performance. This aligns with prior findings indicating that smaller models are less capable of providing high-quality feedback (Saunders et al. 2022; Bai et al. 2022).
Discussion
Regarding RQ1, the results demonstrate that ProRefine is a broadly applicable method that utilizes textual feedback to improve LLM performance at inference time. The “performance gap bridging” effect is particularly noteworthy, suggesting that ProRefine may serve as an effective alternative to simply scaling up model size, potentially avoiding costly fine-tuning an advantage in resource-constrained settings.
@@ Discussion
Regarding RQ3, the results highlight that employing a high-quality verifier is crucial for significantly improving task performance at inference time. We observe some cases where “no verifier” outperforms the “verifier” setting, which indicates the verifier incorrectly accepted a flawed initial answer, thereby preventing the refinement process from correcting the error. This reveals a trade-off: the verifier reduces computational cost on correct answers but risks prematurely halting on incorrect ones. The superior results of the “optimal verifier” highlight the critical role of verifier accuracy. Beyond enhancing performance, the verifier also reduces computational cost during inference by guiding the refinement process. Moreover, it opens up promising avenues for future work, where an optimizer could be designed to maximize rewards guided by the verifier’s assessments. ProRefine can offer a degree of interpretability by exposing the outputs from $LLM_{feedback}$, allowing insights into the model’s reasoning process. Figures 2 and 5 demonstrate cases where model feedback successfully improves the output and where it fails, respectively. Although evaluated on reasoning and math tasks, ProRefine is general and applicable to other tasks.
Figure 5: This figure is an instance when $LLM_{optimizer}$ is not aligned with the feedback from $LLM_{feedback}$ and misses important guiding steps. The framework is similar to Figure 2 when $LLM_{optimizer}$ conveys feedback effectively. We’ve observed a few failed instances following this pattern.
A crucial consideration for ProRefine is the trade-off between its accuracy gains and the increased computational cost at inference time. Each refinement step requires additional calls to the $LLM_{feedback}$ and $LLM_{optimizer}$, making any single query more expensive to process than a standard single-pass generation. However, this per-query cost should be evaluated within ProRefine’s intended hybrid-model deployment. The strategy is not to run refinement on every query, but to use it as an on-demand intervention precisely when a more efficient base model fails. Therefore, the overall system cost is not a simple sum of expensive refinement processes. Instead, it is a blend of low-cost successes from the base model and high-value, targeted corrections. Moreover, the cost is still considerably lower than full model retraining or fine-tuning. Our results support this approach’s practicality: Figure 4 shows that the average number of refinement iterations is typically low, ensuring the per-incident cost of intervention is contained. This cost-accuracy balance can be further optimized by tuning hyperparameters like feedback granularity ($k$) and maximum iterations ($n$).
Limitations and Future Work
@@
  • Computational Cost and Practicality: While ProRefine is designed for cost-effective hybrid deployments, its iterative process inherently increases inference-time latency and computational cost compared to a single-pass query. The cost-benefit of this trade-off must be carefully evaluated for each specific application, as its viability depends on the base model’s failure rate and the relative costs of the LLMs involved.
  • Generalizability: Our evaluation is currently focused on mathematical and multi-step reasoning tasks. Further research is needed to assess performance across a broader range of reasoning tasks and domains. Our method is also sensitive to hyperparameters and requires manual tuning. Developing more robust, automated, or adaptive methods for setting parameters would enhance the method’s usability.
  • Dependence on High-Quality Feedback: The system’s performance is dependent on the quality of the $LLM_{feedback}$. Future work could explore using a specialized “critic” model or fine-tuning feedback models to improve diagnostic accuracy. Furthermore, using LLMs for evaluation introduces potential biases and more comprehensive human evaluations and robust methods are need for mitigating evaluator bias.
  • Stability of the Refinement Loop: The iterative nature of ProRefine lacks a formal convergence guarantee. In some cases, the refinement process can suffer from prompt degradation after many iterations or plateau before reaching an optimal solution. Investigating methods to ensure stable and monotonic improvement is a key area for future research.
Conclusion
We introduced ProRefine, a novel, practical, and inference-time prompt optimization method for agentic workflows. ProRefine leverages LLM-generated textual feedback to dynamically refine prompts, leading to significant performance improvements on multi-step reasoning tasks without requiring additional training or ground-truth labels. Our results demonstrate its ability to bridge the performance gap between smaller and larger LLMs, making it a key enabler for more efficient and cost-effective hybrid-model deployments. The inference-time nature of ProRefine makes it readily deployable for on-demand reasoning correction, contributing to more adaptable and accessible AI systems. Future work will explore applying this framework to new domains, developing more sophisticated feedback and optimizer agents, and exploring adaptive policies for hyperparameter tuning to further optimize the cost-performance trade-off.